# Manvi Jha Electrical and Computer Engineering

### v1

@@ Abstract
Index Terms:
Large Language Models (LLMs), Microprocessor, Verilog
@@ I Introduction
Motivated by this insight, we introduced PREFACE [6], our lightweight and scalable RL framework in which a Small Language Model (SLM) is trained to generate optimized prompts for a fixed, general-purpose LLM, allowing accurate, formally verified code synthesis without costly fine-tuning of the base model. Building on PREFACE, we then present Proof2Silicon, which embeds this same RL-driven prompt adaptation mechanism into a full-stack high-level synthesis pipeline: it takes the verified code output from the LLM and automatically transforms it into synthesizable HLS C (via Dafny’s Python backend and PyLog [7]) and ultimately into RTL through Vivado HLS. The details of PREFACE and its integration into the Proof2Silicon workflow are described in Sectionss III and IV.
@@ II-B Dafny Generation & Verification with LLMs
Figure 1: Overview of the PREFACE [6] framework-forming the prompt-optimization core embedded within the Proof2Silicon pipeline.
@@ III-A Pipeline Architecture
PREFACE [6] is an RL-guided pipeline developed to synthesize formally verified Dafny code, a language designed for rigorous specification and automated verification. It centers on an RL-driven Small Language Model (SLM) that refines the prompts to a frozen LLM $\mathcal{M}$ for the synthesis of Dafny code. Given task specifications $\tau$ (pre / postconditions), a handcrafted prompt $p_{0}=\varphi(\tau)$ is submitted to $\mathcal{M}$, producing a candidate code $c_{t}$. The Dafny verifier then computes the error count $e_{t}$ and feedback $o_{t}$. On success ($e_{t}=0$), the loop terminates; otherwise, the SLM ingests $(p_{t},c_{t},o_{t})$ to sample a prompt edit $\Delta p_{t}$, yielding $p_{t+1}=p_{t}\oplus\Delta p_{t}$. This cycle repeats for up to $T_{\max}$ iterations.
@@ III-B MDP Formulation
  • Action $a_{t}$ is a discrete, token-level modification $\Delta p_{t}$ drawn from the SLM’s vocabulary. $(a_{t}\in\mathcal{A})$
@@
  • $\gamma=0.99$ emphasizes long-term verification success.
@@ III-C Policy Optimization via PPO
while the critic minimizes $L_{\mathrm{critic}}(\phi)=\mathbb{E}_{t}\bigl{[}(V_{\phi}(s_{t})-G_{t})^{2}\bigr{]}$. We use Adam with gradient-norm clipping for stable convergence. By restricting training to the lightweight SLM, PREFACE achieves formally verified code synthesis without fine-tuning the base LLM, forming the core of Proof2Silicon.
Figure 3: Proof2Silicon end-to-end workflow: an RL-driven Small Language Model (SLM) iteratively refines meta-prompts to a frozen LLM under guidance from the integrated Dafny verifier, producing formally verified Dafny code. Verified code is then compiled to Python, transformed into hardware-aware C via the PyLog backend, and synthesized to RTL with Vivado HLS on a Zynq-7000 SoC.
@@ IV-E FPGA Hardware Synthesis and Verification
The final stage uses Xilinx Vivado HLS to synthesize the generated C kernel into RTL, producing detailed reports on timing (clock-period estimates, initiation interval) and resource utilization. Although we do not perform functional co-simulation, this step verifies that the design is structurally FPGA-compatible and meets target performance and area constraints.
In this way, Proof2Silicon seamlessly bridges formally-verified Dafny specifications with practical hardware synthesis—automatically generating synthesizable RTL from natural language descriptions while providing confidence in timing and resource metrics.
Figure 3 illustrates the end-to-end Proof2Silicon workflow. First, the embedded PREFACE framework uses a trained Small Language Model (SLM) to generate meta-prompts that guide a frozen LLM in producing candidate Dafny programs. Each program is immediately checked by an integrated Dafny verifier, and only those that pass formal verification proceed. Verified Dafny code is then compiled to Python using Dafny’s compiler, yielding code that still contains Dafny-specific constructs. This Python code is then reformatted via our automation script to be converted to a PyLog-friendly script. It is then fed into our PyLog transformation pipeline, which rewrites it into HLS C kernels. Finally, these kernels are synthesized with Vivado HLS (10 ns target clock) on the xc7z020clg484-1 device, producing RTL IP cores along with detailed timing and resource utilization reports. By coupling RL-driven prompt optimization with formal verification and automated high-level synthesis, Proof2Silicon delivers correctness-by-construction hardware directly from high-level algorithmic descriptions.
@@ V Results and Evaluation
In this section, we provide an end-to-end evaluation of the Proof2Silicon workflow, beginning with PREFACE’s formally verified Dafny code synthesis results and concluding with the synthesis readiness and performance metrics of the generated HLS C codes. We evaluated the performance of several state-of-the-art language models in a representative subset of 100 tasks sampled from DafnyBench [8]. The models tested include ChatGPT-4o [26], ChatGPT-o1 [27], Qwen2.5-Coder-14B [28], Qwen2.5-7B [29] and Gemini-2-Flash.
@@ V-A Embedded PREFACE results
TABLE I: Verification success rates on a 100-task benchmark for various LLMs using prompts through PREFACE pipeline from both untrained and  100-hr-trained small language model, comparing single-shot (W/O Feedback) versus iterative error-feedback (With Feedback) modes. We employ the "One-line and Detailed Description" prompting mechanism; values are color-coded green for improvements and red for degradations relative to their respective baseline values.
@@
PREFACE was trained for approximately 100 h on two NVIDIA H100 GPUs, during which shaped rewards, policy loss, and value loss were recorded over 2000+ episodes (Figure 4). Both policy and value losses steadily declined, while episode rewards rose despite early-stage penalties, indicating effective convergence and improved prompt optimization.
On a 100-task DafnyBench subset (Table I), static prompting baselines were outperformed by SLM-tuned prompting even without feedback, achieving 23% for both ChatGPT-4o and ChatGPT-o1, 16% for Qwen2.5-Coder-14B, 3% for Qwen2.5-7B, and 29% for Gemini-2-Flash. Enabling the full verifier feedback loop further boosted these rates to 50%, 52%, 31%, 11%, and 55%, respectively, demonstrating PREFACE’s consistent, model-agnostic enhancements in formal verification success.
@@ V-B HLS Synthesis and Performance Results
As shown in Table II, Proof2Silicon achieves robust synthesis readiness across diverse LLMs and feedback modes. Without verifier feedback, the pipeline synthesizes 52.1% of ChatGPT-o1-mini’s verified codes and up to 72.4% for Gemini-2-Flash, with mean synthesis times of 32.8–33.9 s and peak memory under 705 MB. Enabling the full verifier feedback loop further improves synthesis rates—ChatGPT-o1-mini rises to 55.8%, while synthesis latencies remain stable (32.8–34.2 s) and memory usage exhibits minimal variance (678–683 MB). These results demonstrate that embedding PREFACE’s prompt-optimization core within a hardware synthesis pipeline yields high end-to-end success: the majority of formally verified algorithms not only translate cleanly to HLS C but also compile and synthesize into FPGA-ready RTL with predictable performance and resource consumption. Such consistency underlines Proof2Silicon’s potential for automated, correctness-driven hardware generation in safety-critical and high-throughput applications.

### v2

# Manvi Jha    Jiaxin Wan    Deming Chen

@@ I Introduction
Motivated by this insight, we introduced PREFACE [6], our lightweight and scalable RL framework in which a Small Language Model (SLM) is trained to generate optimized prompts for a fixed, general-purpose LLM, allowing accurate, formally verified code synthesis without costly fine-tuning of the base model. Building on PREFACE, we then present Proof2Silicon, which embeds this same RL-driven prompt adaptation mechanism into a full‑stack high‑level synthesis pipeline: it takes the verified code output from the LLM and automatically transforms it into synthesizable HLS C (via Dafny’s Python backend and PyLog [7]) and ultimately into RTL through Vivado HLS. The details of PREFACE and its integration into the Proof2Silicon workflow are described in Sections III and IV.
@@ II-B Dafny Generation & Verification with LLMs
Figure 1: Overview of the PREFACE [6] framework-forming the prompt‑optimization core embedded within the Proof2Silicon pipeline.
@@ III-A Pipeline Architecture
PREFACE [6] is an RL-guided pipeline developed to synthesize formally verified Dafny code, a language designed for rigorous specification and automated verification. It centers on an RL‑driven Small Language Model (SLM) that refines the prompts to a frozen LLM $\mathcal{M}$ for the synthesis of Dafny code. Given task specifications $\tau$ (pre / postconditions), a handcrafted prompt $p_{0}=\varphi(\tau)$ is submitted to $\mathcal{M}$, producing a candidate code $c_{t}$. The Dafny verifier then computes the error count $e_{t}$ and feedback $o_{t}$. On success ($e_{t}=0$), the loop terminates; otherwise, the SLM ingests $(p_{t},c_{t},o_{t})$ to sample a prompt edit $\Delta p_{t}$, yielding $p_{t+1}=p_{t}\oplus\Delta p_{t}$. This cycle repeats for up to $T_{\max}$ iterations.
@@ III-B MDP Formulation
  • Action $a_{t}$ is a discrete, token‑level modification $\Delta p_{t}$ drawn from the SLM’s vocabulary. $(a_{t}\in\mathcal{A})$
@@
  • $\gamma=0.99$ emphasizes long‑term verification success.
@@ III-C Policy Optimization via PPO
while the critic minimizes $L_{\mathrm{critic}}(\phi)=\mathbb{E}_{t}\bigl[(V_{\phi}(s_{t})-G_{t})^{2}\bigr]$. We use Adam with gradient‑norm clipping for stable convergence. By restricting training to the lightweight SLM, PREFACE achieves formally verified code synthesis without fine‑tuning the base LLM, forming the core of Proof2Silicon.
Figure 3: Proof2Silicon end‑to‑end workflow: an RL‑driven Small Language Model (SLM) iteratively refines meta‑prompts to a frozen LLM under guidance from the integrated Dafny verifier, producing formally verified Dafny code. Verified code is then compiled to Python, transformed into hardware‑aware C via the PyLog backend, and synthesized to RTL with Vivado HLS on a Zynq‑7000 SoC.
@@ IV-E FPGA Hardware Synthesis and Verification
The final stage uses Xilinx Vivado HLS to synthesize the generated C kernel into RTL, producing detailed reports on timing (clock‑period estimates, initiation interval) and resource utilization. Although we do not perform functional co‑simulation, this step verifies that the design is structurally FPGA‑compatible and meets target performance and area constraints.
In this way, Proof2Silicon seamlessly bridges formally‑verified Dafny specifications with practical hardware synthesis—automatically generating synthesizable RTL from natural language descriptions while providing confidence in timing and resource metrics.
Figure 3 illustrates the end‑to‑end Proof2Silicon workflow. First, the embedded PREFACE framework uses a trained Small Language Model (SLM) to generate meta‑prompts that guide a frozen LLM in producing candidate Dafny programs. Each program is immediately checked by an integrated Dafny verifier, and only those that pass formal verification proceed. Verified Dafny code is then compiled to Python using Dafny’s compiler, yielding code that still contains Dafny‑specific constructs. This Python code is then reformatted via our automation script to be converted to a PyLog-friendly script. It is then fed into our PyLog transformation pipeline, which rewrites it into HLS C kernels. Finally, these kernels are synthesized with Vivado HLS (10 ns target clock) on the xc7z020clg484‑1 device, producing RTL IP cores along with detailed timing and resource utilization reports. By coupling RL‑driven prompt optimization with formal verification and automated high‑level synthesis, Proof2Silicon delivers correctness‑by‑construction hardware directly from high‑level algorithmic descriptions.
@@ V Results and Evaluation
In this section, we provide an end‑to‑end evaluation of the Proof2Silicon workflow, beginning with PREFACE’s formally verified Dafny code synthesis results and concluding with the synthesis readiness and performance metrics of the generated HLS C codes. We evaluated the performance of several state-of-the-art language models in a representative subset of 100 tasks sampled from DafnyBench [8]. The models tested include ChatGPT-4o [26], ChatGPT-o1 [27], Qwen2.5-Coder-14B [28], Qwen2.5-7B [29] and Gemini-2-Flash.
@@ V-A Embedded PREFACE results
TABLE I: Verification success rates on a 100‐task benchmark for various LLMs using prompts through PREFACE pipeline from both untrained and  100-hr-trained small language model, comparing single‐shot (W/O Feedback) versus iterative error‐feedback (With Feedback) modes. We employ the "One-line and Detailed Description" prompting mechanism; values are color‐coded green for improvements and red for degradations relative to their respective baseline values.
@@
PREFACE was trained for approximately 100 h on two NVIDIA H100 GPUs, during which shaped rewards, policy loss, and value loss were recorded over 2000+ episodes (Figure 4). Both policy and value losses steadily declined, while episode rewards rose despite early‐stage penalties, indicating effective convergence and improved prompt optimization.
On a 100‐task DafnyBench subset (Table I), static prompting baselines were outperformed by SLM‑tuned prompting even without feedback, achieving 23% for both ChatGPT‑4o and ChatGPT‑o1, 16% for Qwen2.5‑Coder‑14B, 3% for Qwen2.5‑7B, and 29% for Gemini‑2‑Flash. Enabling the full verifier feedback loop further boosted these rates to 50%, 52%, 31%, 11%, and 55%, respectively, demonstrating PREFACE’s consistent, model‑agnostic enhancements in formal verification success.
@@ V-B HLS Synthesis and Performance Results
As shown in Table II, Proof2Silicon achieves robust synthesis readiness across diverse LLMs and feedback modes. Without verifier feedback, the pipeline synthesizes 52.1% of ChatGPT‑o1-mini’s verified codes and up to 72.4% for Gemini‑2‑Flash, with mean synthesis times of 32.8–33.9 s and peak memory under 705 MB. Enabling the full verifier feedback loop further improves synthesis rates—ChatGPT‑o1-mini rises to 55.8%, while synthesis latencies remain stable (32.8–34.2 s) and memory usage exhibits minimal variance (678–683 MB). These results demonstrate that embedding PREFACE’s prompt‑optimization core within a hardware synthesis pipeline yields high end‑to‑end success: the majority of formally verified algorithms not only translate cleanly to HLS C but also compile and synthesize into FPGA‑ready RTL with predictable performance and resource consumption. Such consistency underlines Proof2Silicon’s potential for automated, correctness‑driven hardware generation in safety‑critical and high‑throughput applications.